# Are Colors Quanta of Light for Human Vision? A Quantum Cognition Study of Visual Perception

### v1

@@ 2 Categorical Perception, Quantization and Colors
The color system used both by Berlin and Kay and by Rosch, and in which they thus define the eleven prototype colors belonging to the English language, is the Munsell color system, a three-dimensional system where the sizes of the three parameters, hue, chroma and value define each color. Berlin and Kay use their original nomenclature of ‘focus colors’ for the prototype colors, and Rosch also uses this nomenclature in her earliest articles on colors (Rosch Heider, 1971, 1972; Mervis et al., 1975). Since it was mainly her research on colors that made Rosch develop the prototype theory for concepts, she calls these focus colors in her later work the prototype colors and we will use this terminology in our writing on colors.
@@
In the next section we examine the connection between categorical perception and quantum measurement, more specifically, we look in detail at the situation of a two-dimensional quantum entity and show how its measurement dynamics is the underlying ground for the warping mechanism at work in the phenomenon of categorical perception.
3 Quantum Measurement and the Bloch Model
We will work with the Bloch representation of a two-dimensional quantum entity that we will analyze with the intention of identifying the mechanism of categorical perception at work. Traditionally, the Bloch model is only a representation of the set of quantum states, pure states and density states, also called mixed states, of a two-dimensional quantum entity, however, in the 1980s an extension of the Bloch model was worked out in which also the measurements could explicitly be represented (Aerts, 1986). In later years the original extension was elaborated and perfected (Aerts & Sassoli de Bianchi, 2014, 2016), and since this extension contains also a specification of the change provoked by the measuring apparatus on the state of the measured upon entity, we will use it for our analysis, because it will more easily allow us to identify the mechanism of categorical perception. We introduce the extended Bloch model for a two-dimensional quantum entity step by step, explaining details as they arise, the intention being to make the content of our article accessible to readers who are not experts in quantum physics. In parallel, we also introduce the elements of the traditional complex Hilbert space formalism of a two-dimensional quantum entity. That we speak of a ‘two-dimensional’ quantum entity, by the way, refers to that its Hilbert space has dimension 2. Within the discipline of quantum computation a two-dimensional quantum entity is called a ‘qubit’.
A two-dimensional complex Hilbert space $\mathcal{H}$ is a vector space of dimension 2 over the complex numbers with an inner product, which is a map, conjugate linear in its first variable and linear in its second, to the set ${\mathbb{C}}$ of complex numbers
$\langle\ |\ \rangle:\mathcal{H}\times\mathcal{H}\rightarrow{\mathbb{C}}$ (1)
$\langle a\psi_{1}+b\psi_{2}|\psi\rangle=a^{*}\langle\psi_{1}|\psi\rangle+b^{*}\langle\psi_{2}|\psi\rangle$ (2)
$\langle\psi|a\psi_{1}+b\psi_{2}\rangle=a\langle\psi|\psi_{1}\rangle+b\langle\psi|\psi_{2}\rangle$ (3)
$\langle\psi_{1}|\psi_{2}\rangle=\langle\psi_{2}|\psi_{1}\rangle^{*}$ (4)
The inner product allows to express when two states are orthogonal to each other, namely, if their inner product is equal to zero. It also introduces a norm on the Hilbert space, the inner product of a vector with itself is the square of the length of this vector.
In Figures 1 and 2 we consider the basis structure of the Bloch model for a two-dimensional quantum entity, namely the Bloch sphere, which is a sphere in three-dimensional Euclidean space with radius equal to 1. The points of the surface of this sphere represent the pure states of the considered quantum entity. These are the states that the entity can occupy independently of being measured upon or not, thus they describe the reality of the considered entity. In Hilbert space, these pure states are represented by unit vectors of the Hilbert space. We promised that we wanted to keep the content of this article understandable for those who do not know the mathematical technicalities of the quantum formalism, and we will do our best to keep this promise. Nevertheless, we are obliged to introduce some mathematical elements that are typical of the quantum formalism. The pure states of two dimensional quantum entity, hence a qubit, or the spin of a quantum particle with half integer spin, are represented in the extended Bloch model by the points of the surface of the Bloch sphere.
Figure 1: A representation of our extended Bloch model for a two-dimensional quantum entity. A little ball is placed in a point $A$ on the surface of a sphere and this represents the state of the entity in $A$ before the measurement. On a central axis of the sphere between the points $A_{down}$ and $A_{up}$ situated on the surface of the sphere is placed an elastic and a first phase of the measurement consists of the ball falling through the interior of the sphere orthogonally to the elastic and sticking to it in a point $A^{\prime}$ on the centerline between $A_{down}$ and $A_{up}$. The second phase of the measurement consists of the elastic breaking in one of its points randomly and uniformly, hence such that the probability of breaking in a piece of the elastic is proportional to the length of this piece. A a consequence the ball is moved upwards ending up in the point $A_{up}$ or downwards ending up in the point $A_{down}$ depending on whether the elastic breaks in the part down to $A^{\prime}$ or the part up to $A^{\prime}$.
@@ 3 Quantum Measurement and the Bloch Model
We will denote these points by vectors in the three-dimensional Euclidean space, indeed, the Bloch sphere is contained in this three-dimensional Euclidean space, by the so-called spherical coordinates $r$, $\theta$, and $\phi$. The distance from the center of the Bloch sphere to a point on the surface of the Bloch sphere is traditionally denoted in spherical coordinates by $r$, and called the ‘radial distance’. Since our Bloch sphere has radius equal to 1, for any point on the surface we have $r=1$. The angle that the line connecting the center of the Bloch sphere with the considered point of the surface of the Bloch sphere with the axis of the Bloch sphere is denoted by $\theta$, and it thus varies between values 00 and $\pi$, it is equal to 00 if the considered point coincides with the North Pole of the Bloch sphere and is called the polar angle (see Figure 2). Hence, it is equal to $\pi$ if it coincides with the South Pole  [^2] We denote the magnitude of angles by the unit of a radian, which makes 180 degrees equal to $\pi$ radians. A second angle is still needed so that the values of $r$, and $\theta$, and the value of that second angle, traditionally denoted $\phi$ and called the azimuthal angle, would uniquely determine each point of the three dimensional Euclidean space. This angle $\phi$ is chosen from a vertical plane through the axis of the Bloch sphere rotating around that axis reaching the other points not in the vertical plane, for the points in the plane $\phi$ is equal to zero or $2\pi$. This means that for values of $\phi$ varying between 00 and $2\pi$, all points of three-dimensional Euclidean space can be reached and hence the spherical coordinates ($r$, $\theta$, $\phi$) form complete coordination of the three dimensional Euclidean space. For $r=1$, by varying the values of $\theta$ and $\phi$, all points of the surface of the Bloch sphere are reached (see Figures 1 and 2). So the point $A$ has spherical coordinates $(1,\theta,\phi)$, and cartesian coordinates $(x,y,z)$, where
$x$ $=$ $\sin\theta\sin\phi$ (5)
$y$ $=$ $\sin\theta\cos\phi$ (6)
$z$ $=$ $\cos\theta$ (7)
for a cartesian coordinate system with its origin in the center of the sphere.
The coordinates, be they spherical $(r,\theta,\phi)$ or cartesian $(x,y,z)$, that we have considered so far, describe a three-dimensional Euclidean space, of which the Bloch sphere is a part.
Figure 2: A three dimensional representation of the Extended Bloch Model. The little ball is in point $A$ with spherical coordinates $(\rho,\theta,\phi)$ and Hilbert space density operator coordinates $D_{(\rho,\theta,\phi)}$. The process of measurement proceeds like described in detail in Figure 1 by the ball falling orthogonally to the elastic and sticking to it in point $A^{\prime}$. Then the elastic breaks uniformly in one of its points an pulls the ball upwards to end in $A_{up}$ or downwards to end in $A_{down}$.
However, we have explicitly discussed a two-dimensional quantum entity that we want to describe using its Bloch representation, while the Cartesian and spherical coordinates describe a three-dimensional space. If one remembers the representation of the complex numbers in the Euclidean plane, a mathematics subject of high school education, then it comes as no surprise that a representation of a two-dimensional complex vector space needs three dimensions to be represented in a Euclidean real space. Indeed, the set of complex numbers, if one wishes to think of it that way, is a one-dimensional complex vector space and it needs a plane, the so called complex plane, hence a two-dimensional space, if represented by points with coordinates that are real numbers, such as is the case by polar coordinates in the complex plane. Let us introduce a general vector from the two-dimensional vector space, and we immediately refer to the spherical coordinates to indicate to which point of the Bloch sphere that vector belongs.
$|\theta,\phi\rangle=(\cos{\theta\over 2}e^{-i{\phi\over 2}},\sin{\theta\over 2}e^{i{\phi\over 2}})$ (8)
We do not yet directly explain here why precisely this vector $|\theta,\phi\rangle$ of the two-dimensional complex Hilbert space is the one we let correspond to that point $(1,\theta,\phi)$ of the three-dimensional Euclidean space, the analysis we make in the course of this article will clarify this.
The notation $|\theta,\phi\rangle$ is the one introduced by Paul Dirac, one of the founding fathers of quantum mechanics, along with Werner Heisenberg and Erwin Schrödinger (Dirac, 1939). Dirac considered the inner product as a fundamental operational element, the bracket, $\langle\theta,\phi|\theta,\phi\rangle$, and called the right side of this bracket, the ket, $|\theta,\phi\rangle$ and the left side, the bra $\langle\theta,\phi|$. Hence the notation $|\theta,\phi\rangle$ as the right side, thus as ket, was chosen as a representation of the state of the considered entity by Dirac because the inner product is taken to be ‘linear’ in the second variable while ‘conjugate linear’ in the first variable  [^3] This choice, by the way, is sometimes not followed by mathematicians who define the inner product in a complex Hilbert space as linear in the first variable and conjugate linear in the second.. In matrix notation, in the case of a two dimensional quantum entity, the ket is written as a ‘one column’ to ‘two row’ matrix, and the bra as a ‘one row’ to ‘two column’ matrix, plus a complex conjugation of the terms of the matrix for the bra.
$|\theta,\phi\rangle$ $=$ $\begin{pmatrix}\cos{\theta\over 2}e^{-i{\phi\over 2}}\\ \sin{\theta\over 2}e^{i{\phi\over 2}}\end{pmatrix}$ (9)
$\langle\theta,\phi|$ $=$ $\begin{pmatrix}\cos{\theta\over 2}e^{i{\phi\over 2}}&\sin{\theta\over 2}e^{-i{\phi\over 2}}\end{pmatrix}$ (10)
We thus find by a direct matrix calculation that the ket is normalized as a state, indeed we have
$\langle\theta,\phi|\theta,\phi\rangle$ $=$ $\begin{pmatrix}\cos{\theta\over 2}e^{i{\phi\over 2}}&\sin{\theta\over 2}e^{-i{\phi\over 2}}\end{pmatrix}\begin{pmatrix}\cos{\theta\over 2}e^{-i{\phi\over 2}}\\ \sin{\theta\over 2}e^{i{\phi\over 2}}\end{pmatrix}$ (11)
$=$ $\cos^{2}{\theta\over 2}+\sin^{2}{\theta\over 2}=1$
Dirac also introduced in his bra-ket calculus the possibility of multiplying the ket by the bra in the reverse order of the bra-ket, that is, as ket-bra. Whereas the bra-ket multiplication always results in a complex number, more specifically the number 1 in the example we considered here, the reverse multiplication $|\theta,\phi\rangle\langle\theta,\phi|$ leads to a more complex quantity. We can calculate this more complex quantity directly in the case of the two-dimensional quantum entity we are considering, and by the representations of ket and bra by the two-by-one matrices, as we just explained. We then get
$|\theta,\phi\rangle\langle\theta,\phi|$ $=$ $\begin{pmatrix}\cos{\theta\over 2}e^{-i{\phi\over 2}}\\ \sin{\theta\over 2}e^{i{\phi\over 2}}\end{pmatrix}\begin{pmatrix}\cos{\theta\over 2}e^{i{\phi\over 2}}&\sin{\theta\over 2}e^{-i{\phi\over 2}}\end{pmatrix}$ (12)
$=$ $\begin{pmatrix}\cos^{2}{\theta\over 2}&\cos{\theta\over 2}\sin{\theta\over 2}e^{-i\phi}\\ \cos{\theta\over 2}\sin{\theta\over 2}e^{i\phi}&\sin^{2}{\theta\over 2}\end{pmatrix}$
a ‘two-by-two’ matrix. This two-by-two matrix represents an operator within the quantum Hilbert space formalism called an orthogonal projection, but before we continue on that, we wish to specify using the extended Bloch model what happens during a quantum measurement.
Within the extended Bloch model, the measurement is represented by a piece of elastic stretched on the centerline between two diametrically opposed points of the sphere, $A_{down}$ and $A_{up}$ (see Figures 1 and 2). On the surface of the sphere is a small ball that sticks to the sphere at a point $A$. This ball in a point represents the state in which the quantum entity is. The measurement then occurs as follows. First, the ball falls orthogonally on the elastic and sticks to it at the point $A^{\prime}$ where this orthogonal fall brings it. Then the elastic breaks randomly in one of its points. If the breaking point is below the point $A^{\prime}$, then the unbroken part of the elastic pulls the ball up so that it ends up in the point $A_{up}$. However, if the breaking point is above the point $A^{\prime}$, then the ball is pulled down by the unbroken part of the elastic, and ends up in the point $A_{down}$. Thus, the whole of the measurement results in the ball from the point $A$ ending up in one of two points $A_{up}$ or $A_{down}$.
Now suppose that the elastic possesses a uniform breaking pattern, which means that the point at which the elastic breaks randomly will lie in a certain interval of the elastic with a probability proportional to the length of that interval. We can then calculate the probabilities $P(A\mapsto A_{up})$ and $P(A\mapsto A_{down})$ with which the measurement will carry the ball from $A$ to $A_{up}$ or to $A_{down}$, and in this way will change the state of the quantum entity, from the simple geometry of the configuration. Note indeed that we have that the length $L_{2}$ of the elastic under the point $A^{\prime}$ equals $1+\cos\theta$ and the length $L_{1}$ of the elastic above the point $A^{\prime}$ equals $1-\cos\theta$, which hence, taking into account the uniform nature of the breaking pattern of the elastic, gives us
$P(A\mapsto A_{up})$ $=$ ${1+\cos\theta\over 2}=\cos^{2}{\theta\over 2}$ (13)
$P(A\mapsto A_{down})$ $=$ ${1-\cos\theta\over 2}=\sin^{2}{\theta\over 2}$ (14)
The probabilities (13) and (14) match the quantum probabilities of the spin of a spin 1/2 quantum particle or a qubit. The ball in point $A$ is then in a state for the spin of an angle $\theta$ with the $z$-axis, and the elastic lies on this $z$-axis.
The part of this quantum measurement that takes place at the so-called pure states is described in a complete way in the quantum Hilbert space formalism. The complex Hilbert space in this case is two-dimensional, with basis vectors $|0,0\rangle$ and $|\pi,0\rangle$, which in the Bloch representation correspond to points on the Bloch sphere located at the North and South Poles, respectively. In Hilbert space these basis vectors, and we use now their notation as one column and two row matrices, are
$|0,0\rangle=\begin{pmatrix}\cos{\theta\over 2}e^{-i{\phi\over 2}}\\ \sin{\theta\over 2}e^{i{\phi\over 2}}\end{pmatrix}_{(\theta=0,\phi=0)}=\begin{pmatrix}1\\ 0\end{pmatrix}$ (15)
$|\pi,0\rangle=\begin{pmatrix}\cos{\theta\over 2}e^{-i{\phi\over 2}}\\ \sin{\theta\over 2}e^{i{\phi\over 2}}\end{pmatrix}_{(\theta=\pi,\phi=0)}=\begin{pmatrix}0\\ 1\end{pmatrix}$ (16)
A general pure state $|\theta,\phi\rangle$, in this Hilbert space formalism, is a normalized linear combination, called superposition in the quantum jargon, of these basis vectors
$|\theta,\phi\rangle=\begin{pmatrix}\cos{\theta\over 2}e^{-i{\phi\over 2}}\\ \sin{\theta\over 2}e^{i{\phi\over 2}}\end{pmatrix}=a\begin{pmatrix}1\\ 0\end{pmatrix}+b\begin{pmatrix}0\\ 1\end{pmatrix}$ (17)
${\rm with}\ a=\cos{\theta\over 2}e^{-i{\phi\over 2}}{\rm and}\ b=\sin{\theta\over 2}e^{i{\phi\over 2}}{\rm and}\ |a|^{2}+|b|^{2}=1$ (18)
with superposition coefficients $a$ and $b$.
We now know the quantum probabilities, namely $\cos^{2}{\theta\over 2}$ and $\sin^{2}{\theta\over 2}$, for the state $|\theta,\phi\rangle$ to collapse as a result of a measurement to the state $|0,0\rangle$ or the North Pole of the Bloch sphere, or the state $|0,\pi\rangle$ or the South Pole of the Bloch sphere, respectively. That information gives us a way to determine the mathematical form in the two-dimensional complex Hilbert space that the state $|\theta,\phi\rangle$ possesses. Indeed, lets write the most general mathematical form for this state $|\theta,\phi\rangle$, namely
$|\theta,\phi\rangle=a|0,0\rangle+b|\pi,0\rangle$ (19)
where $a$ and $b$ are two complex numbers such that $|a|^{2}+|b|^{2}=1$, we just apply the superposition principle here, one of the basic principles of quantum theory. And then we must have
$|a|^{2}=\cos^{2}{\theta\over 2}\quad{\rm and}\quad|b|^{2}=\sin^{2}{\theta\over 2}$ (20)
We can solve these equations and should not forget that $a$ and $b$ are complex numbers. A straightforward solution is $a=\cos{\theta\over 2}$ and $b=\sin{\theta\over 2}$, but that is not the general solution. Note that a dependence of a solution on $\phi$ must vanish in $|a|^{2}$ and $|b|^{2}$, which means that this dependence on $\phi$ is in the form of a phase factor $e^{i\phi}$, since such a phase factor disappears if the absolute value of the complex numbers is calculated. This is how we arrive at the general solution $a=\cos{\theta\over 2}e^{-i{\phi\over 2}}$ and $b=\sin{\theta\over 2}e^{i{\phi\over 2}}$, which we introduced in (23), noting that without loss of generality we can multiply with a phase factor such that this specific form arises  [^4] Another choice often found in textbooks is $a=\cos{\theta\over 2}$ and $b=\sin{\theta\over 2}e^{i\phi}$, which equals the solution we use in this article except a phase of $e^{-i{\phi\over 2}}$.. It can be shown that the complex numbers as a number system on which the vector space is built are essential, it is not possible to obtain the probability structure of a qubit with a two-dimensional vector space over the real numbers (Aerts, 1986).
What we are particularly interested in for our present article is the fine structure of the mechanism in the measurement in the extended Bloch model shown in Figures 1 and 2. And before we continue, we wish to note the following. The example depicts a two-dimensional quantum entity and a measurement with two final states $A_{up}$ and $A_{down}$, hence two outcomes. However, in later years, it was shown that a similar model can be built for a $n$-dimensional quantum entity and a measurement with an arbitrary number $n$ of final states, and thus an arbitrary number $n$ of outcomes (Aerts & Sassoli de Bianchi, 2014). We will not explicitly describe these higher dimensional extended Bloch models in the present article, but mention that the simple geometric properties of the configurations lead to exactly the quantum probabilities in a completely similar way than this is the case for this two dimensional quantum entity. The details of the analysis that we will now make for the two dimensional quantum entity can also be made for the higher dimensional quantum entities in a similar way. Of course, that we find exact quantum probabilities in these measurement models also depends on a certain symmetry that we introduced at the measurement level, namely that the elastic breaks with a randomness that is ‘uniformly’ distributed. If the elastic does not possess this symmetry property, we will still find probabilities that do not fit a classical probability model, but they will not be perfect quantum probabilities, e.g. a complex Hilbert space model will not be able to represent them, they will thus be rather quantum-like. More general quantum formalism than this of standard quantum mechanics in a complex Hilbert space, and not based on vectors representing states, can then be used (Aerts, 1982, 1992; Gudder & Zanghi, 1984).

### v2

@@ 2 Categorical Perception, Quantization and Colors
The color system used both by Berlin and Kay and by Rosch, and in which they thus define the eleven prototype colors belonging to the English language, is the Munsell color system, a three-dimensional system where the sizes of the three parameters, hue, chroma and value define each color. Berlin and Kay use their original nomenclature of ‘focus colors’ for the prototype colors, and Rosch also uses this nomenclature in her earliest articles on colors (Rosch Heider, 1971, 1972; Mervis et al., 1975). Since it was mainly her research on colors that made Rosch develop the prototype theory for concepts, she calls these focus colors in her later work the prototype colors, and we will use this terminology in our writing on colors.
@@
In the next section, we discuss the connection between categorical perception and quantum measurement using the results obtained in Aerts et al., (2025). Hence, we look in detail at the situation of a two-dimensional quantum entity and show how its measurement dynamics is the underlying ground for the warping mechanism at work in the phenomenon of categorical perception.
3 Quantum Measurement and Categorical Perception
In this section we will mainly write down what was proved in Aerts et al., (2025), namely that the warping characteristic of the phenomenon of categorical perception is also present in the measurement process of quantum mechanics. The analysis we did in Aerts et al., (2025) is very thorough and detailed, and we will not repeat it here in its entirety. On the other hand, we want to keep the present article self-contained, so we also do not want to mention only the results. More so, we want to reproduce the understanding of ‘why the warping characteristic of the phenomenon of categorical perception is also the warping present in the quantum measurement apparatus’. We can do this in a simplified way because an intermediate result has been proved in Aerts et al., (2025) from which we can start. This allows us to represent the insightful elements of the analysis in Aerts et al., (2025) by using the simple geometry of triangles in a circle in the Bloch representation of a quantum measurement.
In Figure 1, we represent the Bloch sphere of a two-dimensional quantum entity, which is a sphere in three-dimensional Euclidean space with radius equal to 1. The Bloch sphere is the representation of the two-dimensional quantum entity in which we will prove our assertion. A two-dimensional quantum entity is also described in a two-dimensional complex Hilbert space, and in quantum computing it is called a ‘qubit’. Our assertion can also be for any larger dimension, and we will return to this explicitly in future work. Another well-known example of a two-dimensional quantum entity is the spin of a spin 1/2 quantum particle, and we keep this example in mind because it allows us to make some aspects of the proof in this section more demonstrative.
In the Bloch representation, the pure states of the quantum entity are represented by the points on the surface of the Bloch sphere. The density states are represented by the interior points of the Bloch sphere. So, looking at Figure 1, the points $A$, $A_{down}$ and $A_{up}$, all points of the surface of the Bloch sphere, represent pure states of the quantum entity. The point $A^{\prime}$, on the other hand, being an interior point of the Bloch sphere, represents a density state of the quantum entity. These four points also allow us to describe what happens during a measurement in quantum mechanics when this measurement is performed on the two-dimensional quantum entity we are considering. The point $A$ represents the state of the quantum entity before the measurement. For a two-dimensional quantum entity, each measurement always has two possible outcomes and when one of these outcomes takes place, the initial state of the quantum entity, represented by the point $A$ in our case, is changed into one of the two so-called eigenstates of the measurement, in Figure 1 we have represented these eigenstates by the points $A_{down}$ and $A_{up}$. So concretely, after performing the measurement, the quantum entity is either in the pure state $A_{down}$, and we say that the outcome took place which we call ‘down’, or in the pure state $A_{up}$, and then we say that the outcome took place which we call ‘up’.
Figure 1: A three-dimensional representation of the Bloch sphere. The quantum state of the considered quantum entity represented by point $A$ moves as a consequence of decoherence during a measurement on the line orthogonally to the diameter of the sphere to end up in the density state represented by the interior point $A^{\prime}$ of the Bloch sphere and then collapses to one of the two eigen states of the measurement represented by points $A_{down}$ or $A_{up}$ respectively giving outcome ‘down’ or ‘up’.
However, during the quantum measurement, the state of the quantum entity moves along a trajectory in the Bloch sphere. In a first stage, a process of decoherence takes place, the pure state in point $A$ falls orthogonally on the diameter connecting the two eigenstates and thus arrives at the final stage of decoherence in point $A^{\prime}$. We will come back to exactly what this process of decoherence means later. As we already mentioned, the point $A^{\prime}$, being an interior point of the Bloch sphere, represents a density state of the quantum entity. From this point $A^{\prime}$ the so-called quantum collapse takes place, the trajectory becomes discontinuous, and jumps either with outcome ‘down’ to state $A_{down}$, or else with outcome ‘up’ to $A_{up}$, again both pure states, thus the quantum coherence is restored.
@@ 3 Quantum Measurement and Categorical Perception
The angles $\theta$ and $\phi$ are the angles of the so-called spherical coordinates. Specifically, $\theta$ is the angle that the line connecting the center of the Bloch sphere with the considered point $A$ of the surface of the Bloch sphere makes with the axis of the Bloch sphere, and thus varies between values 00 and $\pi$, it is equal to 00 if the point $A$ coincides with the North Pole of the Bloch sphere and it is called the polar angle (see Figure 1). Hence, it is equal to $\pi$ if it coincides with the South Pole  [^2] We denote the magnitude of angles by the unit of a radian, which makes 180 degrees equal to $\pi$ radians. The angle $\phi$ is called the azimuthal angle, it is the angle between a fixed axis in the horizontal plane of the equator and the line that this horizontal plane has in common with the vertical plane through the north and south poles and the point $A$ (see Figure 1). It varies between 00 and $2\pi$.